A firm has total cost of TC(y)=y²+1 and marginal cost of MC(y)=2y. What is the firm's producer's surplus at price $8?
Answer #1

Коли ООО + Producer Svetlis МС изм Total cost ?

The firm's Marginal cost curve is also the firm's supply curve.

So the area below the supply curve or Marginal Cost curve is the sum of Marginal cost from 0 units upto y units which is the Total cost at y units.

Also since the rectangular area depicts the Total Revenue curve.

So, Total Revenue - Total Cost = Producer Surplus

As under equilibrium, P = MC

So, 8 = 2y

y = 8/2 = 4

So, Total Revenue = P*y = 8 * 4 = 32

Total Cost = y​​​​​​2 + 1 (as given above in question)

So, Total Cost = 42 + 1= 16 + 1 = 17

Now, Producer Surplus = 32 - 17 = $15

The firm's producer surplus is $15